On adding agno3 solution into ki solution a negatively charged colloidal sol is obtained when they are in 1. 100ml of 0.1 m agno3+ 100 ml of 0.1m ki

Grade:12th pass

1 Answers

Rituraj Tiwari
3 years ago
In order to form a negatively charged colloidal sol on addingAgNO3​solution into KI solution, KI should be in excess.Among all options, KI is in excess in option B as shown:
50 mL of 0.1 MAgNO3​+ 40 mL of 0.1 M KCl
m moles ofAgNO3​=50×0.1=5
m moles ofKI=50×0.2=10
KI is in excess
So,I−is adsorbed on AgI sol